OMPANYA 49 /35 400 308 2018 17. TRADE & OTHER RECEIVABLES Trade recevable8. Allowance for doubtful debts Other receivables Balance at end of the year 5000 000 44,72038,189 (112)(48) 799 3,803 45,407 39,944 The sverage credit period on sales s 42 days (2015:72 dayst No interest is cherged on trade recevables outstanding Included in the Groups accounits recevable balance are dettorswith a carfying amount of $338,000 201 1,025.000) which are past due at the reporting date but notconsidered doubeful. These rel ate to a number of accounts of which there is no recent history of default The Group hias nat provided for these debtors as there has not been & significant change in credit Quality and the amounts are stit considered recoverable. The ageing of the debtors that are past due but not impaired ars predominanitiy 30 days cor more beyond the dué date of commercial trading terms 2016 $000 2015 $000 17.1 Hovenent in al1owance for doubttu1 debts Balance at beginning ofyear Amouht charged to the statement of comprehensive income Not foreign currency exchange diferences. 48 58 15) 112 46 Balance at end of year ia in determining the tecoverablity of a trade recevable, the Group considers ary change in perceived credit quality of eche trade receivable from the date credit was initially granted up to the repprting date Accordingly, the directars telave th there is no further credit provision required in excess of the allowance for imparment loases

what was the bad debt method that the company used?? How much was their bad debt expense for the year 2016
Answer #1

Bad debt expense for the year 2016 =$69000

Since the opening balance in allowance for doubtful debts at the beginning of year 2016 was $48000. Closing Balance of allowance of doubtful debts are $112000 ($48000+$69000-$5000 for Net difference in foreign currency exchange )

Recording Bad Debts transaction in the Journal- Journal Entry for Bad Debts are mentioned below

Date of transaction Particulars Debit (Dr) Credit(Cr)
Bad Debt Expense A/c Dr 69000
To Allowance for Doubtful Account 69000
Allowance for Doubtful Account Dr 69000
To Accounts Receivable 69000

Bad Debt Expense for the year 2016 is calculated using certain percentage of Net Credit Sales. Since the business environment is highly dynamic, uncertain and unpredictable in nature, most organization create provisions/ allowance for bad & doubtful debts because of conservatism principle.

Calculation of Net Credit Sales for year 2016- It is given to us that average credit period on sale is 42 days

Average collection period = 365/ Accounts Receivable Turnover ratio

42=365/Accounts Receivable Turnover Ratio. Therefore Accounts Receivable Turnover ratio =365/42= 8.7 times

Accounts Receivable Turnover ratio = Net value of Credit Sales/ Average Accounts Receivable during the same time period

Hence Net Credit Sales= Accounts Receivable Turnover ratio * Average Receivables during the same period

Net Credit Sales = 8.7 * (836000+ 1025000)/2)

Net Credit Sales= 8.7*930500 i.e. 8095350

Therefore Bad Debt Method used by the company to calculate provision of bad & doubtful debts are certain percentage of Credit Sales during the year. In the Balance sheet Accounts Receivable will be reduced by $112000 due to the provision for doubtful debts created by the company by following conservatism principle.